OpenStudy (anonymous):

Consider:2(-3i+4)-i(4-i2) Simplify the expression and then determine a and b that produce an equivalent expression of the form a+bi. a. A=8,b=11 b. A=8,b=-11 c. A=7,b=-10 d. A=7,b=9

Nnesha (nnesha):

and start with distributive property \[\rm a(b+c)= a*b+a*c=ab+ac\]